2 yd ³ of soil containing 35% sand was mixed into 6yd ³ of soil containing 15% sand. What is the sand content of the mixture?
Answer by lwsshak3(11628)

2 yd ³ of soil containing 35% sand was mixed into 6yd ³ of soil containing 15% sand. What is the sand content of the mixture?
***
35%*2+15%*6=x%*8
.70+.90=8x
x=1.6/8=.20
sand content of the mixture=20%